The total cost of 8 buckets and 5 mugs is $92, and the total cost of 5 buckets and 8 mugs is $77. Find the cost of one bucket and the cost of one mug.
Answer by checkley77(12844)

8B+5M=92 MULTIPLY BY 8
5B+8M=77 MULTIPLY BY -5 & ADD.
64B+40M=736
-25B-40M=-385
------------------
39B=351
B=351/39
B=$9 COST FOR 1 BUCKET.
8*9+5M=92
72+5M=92
5M=92-72
5M=20
M=$4 COST FOR 1 MUG.
PROOF:
5*9+8*4=77
45+32=77
77=77